OpenStudy (anonymous):

kyla bought 14 yards of green rubber to make a bags for her school she estimates that it takes 2/5 yard to make one bag how many bags can she make? I think 5 and 3/5 because she can make 5 whole bags with 3/5 leftover..

OpenStudy (anonymous):

I like your math, but you can't make 3/5 of a bag

OpenStudy (anonymous):

...so I'd answer 5

OpenStudy (anonymous):

:D right! but wouldnt she have a leftovers?

OpenStudy (anonymous):

one bag takes 2/5 yards. there are totally 14 yards. so total bags = 14/2/5 = 35 bags

OpenStudy (anonymous):

so its, 5 and 3/5 or 35?

OpenStudy (anonymous):

no, d, 14*2 = 28 and 28/5 = 5 and 3/5, not 35

OpenStudy (anonymous):

what it means is, you can make 1 bag with two-fifths of a yard. so with one yard, you can make two and a half bags. so with fourteen yards you can make 14 times 2.5

OpenStudy (anonymous):

agree with dhatraditya

OpenStudy (anonymous):

ok... Ill ask my teacher then! :D

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

the answer is: 5 and 3/5

OpenStudy (anonymous):

can you guys explain how you got 5 and 3/5?